OpenStudy (anonymous):

a decimal is the same when multiplied by 5 as it is when 5 is added to it

OpenStudy (scorcher219396):

Write out an equation based on the information you were given 5x=x+5 -x -x 4x=5 /4 /4 x=5/4 or 1.25 Check yourself: 1.25(5)=6.25 and 1.25+5=6.25 so you're good